OpenStudy (anonymous):

lucy drove from her house to a friends house at an average speed of 40 miles per hour. she returned home along the same route at an average speed of 60 miles per hour. If her total driving time was 1 hour, how many total miles did lucy drive?

OpenStudy (whpalmer4):

She drove 40 mph for \(x\) hours, and 60 mph for \(1-x\) hours. In that time, she drove the same distance: \[40 x = 60(1-x)\]Solve that for \(x\) to find out how long she drove at 40 mph. \(1-x\) gives you the driving time at 60 mph. Multiply speed * time to get distance for each leg.

OpenStudy (anonymous):

so after i multply that, thats the answer?

OpenStudy (whpalmer4):

One step at a time. Find the value of \(x\).

OpenStudy (anonymous):

x=3/5

OpenStudy (whpalmer4):

very good! So she drives for 3/5 of an hour at 40 mph on the trip to her friends house. How far away does her friend live?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

miles away?

OpenStudy (whpalmer4):

Right. And after visiting her friend, she drives back home along the same route. How much total driving does she do? Yes, units are miles because the speeds were in miles/hour and the time was in hours

OpenStudy (anonymous):

60 miles/hour

OpenStudy (whpalmer4):

Let me rephrase my question: "how many miles did she drive?"

OpenStudy (anonymous):

from her friends house to her house?

OpenStudy (whpalmer4):

Reread the problem statement. What are you asked to find?

OpenStudy (whpalmer4):

Essential skill in solving problems: know what you are trying to find!

OpenStudy (anonymous):

trying to find what is the total miles lucy drives

OpenStudy (whpalmer4):

Okay. So, the answer is?

OpenStudy (anonymous):

60 miles

OpenStudy (whpalmer4):

Show me how you arrived at that figure.

OpenStudy (anonymous):

so like at first you get x=3/5 and then 3/5 multiplied by 40= 24 and then 3/5*60=36 so then 36+24=60

OpenStudy (whpalmer4):

No. \(x\) represents the number of hours she drove on the way to her friend's house, traveling at 40 mph. It does NOT represent the number of hours she drove at 60 mph on the return trip along the same route.

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

Stop. Read the problem carefully. Think. Do not use any numbers until you see your mistake.

OpenStudy (anonymous):

idk what i did wrong, idk why i keep getting 60

OpenStudy (whpalmer4):

"She drove 40 mph for x hours, and 60 mph for 1−x hours. "

OpenStudy (whpalmer4):

But even more fundamentally: "she returned home along the same route"

OpenStudy (whpalmer4):

If she takes the same route, isn't the distance going to be the same?!?

OpenStudy (anonymous):

yea but the time isnt the same

OpenStudy (whpalmer4):

BFD. We want to know HOW MANY MILES SHE DROVE

OpenStudy (whpalmer4):

We know that she drove 24 miles on the way to her friend's house. She drove the same route on the way home. How many miles did she drive in total?

OpenStudy (anonymous):

48

OpenStudy (whpalmer4):

YES!!!!

OpenStudy (whpalmer4):

You could also calculate it. \((1-x)\) is the number of hours she spent driving home, at 60 mph. Remember, the total driving time was 1 hour, so if she spent \(x\) driving there, she spent \(1-x\) driving back. \[(1-\frac{3}{5})*60 =\]

OpenStudy (anonymous):

oh, thank you for your help, and im sorry for getting you mad

OpenStudy (whpalmer4):

no, not mad, just frustrated that you couldn't see the end of your nose :-)

OpenStudy (whpalmer4):

Very important to read the problem carefully and understand what you know, and what you need to find. Time spent reading the problem carefully is often more valuable than immediately diving into the calculation...and then when you think you have an answer, read the problem again, and make sure you've answered the right question.

OpenStudy (anonymous):

Oh ok thank you :)